OpenStudy (anonymous):

evaluate the polynomial for x=-2 -4x to the 3rd power + 6x to the 2nd power -7x-1

jhonyy9 (jhonyy9):

\[-4x ^{3}+6x ^{2}-7x-1=-4(-2)^{3}+6(-2)^{2}-7(-2)-1=\]

jhonyy9 (jhonyy9):

=-4(-8)+6(4)+14-1=32+24+14-1=69